# Traditional Methods Outperform Generative LLMs at Forecasting Credit Ratings

### v1

@@ 6. Interpretability
The most obvious advantage of a feature-based system is that important features can be identified. Table 2 provides an example of feature importance that can be used to infer the modality preference of model configurations. It is possible to conduct even more granular feature analysis by looking at the contribution of individual features. Figure 2 shows the partial dependence plot of some of the individual text features on the ”Up” and ”Down” classes. From the plot in Figure 2(a) we can infer that as ratings are discussed more in a company’s filing, there is a reduced chance of the the credit rating being upgraded. We can also infer from Figure 2(b) that as companies talk about receivables - the money owed to the company - there is a reduced chance of the credit rating being downgraded. Both provide valuable insights, and are examples of how a traditional feature-based methodology can be leveraged.
7. Conclusion
@@
The paper shows that while LLMs are good at encoding textual data and inferring signals that traditional methods cannot pick up on, when combined with numerical information in the prompt there is performance deterioration. Traditional methods still outperform LLMs at combining multimodal data. The other advantage is that traditional methods offer increased interpretability and a better understanding of the mechanisms behind certain predictions.
@@
There has been some work jointly encoding text using generative LLMs with time-series information (Liu et al., 2024), but more work needs to be done to determine the best methodology for combining long text sequences with numerical information while utilizing the benefits of generative LLM natural language understanding. This paper shows that combining multimodal information within the prompt is not sufficient.
Appendix A Limitations
@@
The task above uses a balanced dataset, which is good for testing the different methodologies’ ability to discern the signals that are predictive of a rise or fall in credit ratings, but is poor for assessing how good the models would be in a real-world context where almost all of the ratings stay the same. As such, despite the data being taken from across all US equities over a 23-year time-period, the balanced dataset is relatively small, with only 3441 samples for the Lag 1 configuration and 2142 samples for the Lag 4 configuration. There are plenty of prominent datasets that are smaller, but the size reduces the scope for complex and specialised models to be deployed on this task in favour of more robust, simple models.
@@
Another limitation is that the text used in this paper is produced by the company, who have an agenda to convey a positive agenda to investors. More objective publication venues may produce different insights about the future direction of a company.
@@
The LoRA fine-tuning methodology outlined in this paper is a parameter-efficient technique and has been shown to be competitive in a variety of settings (Hu et al., 2021), but can be outperformed in some tasks by full fine-tuning and other adapter-based methods (Xu et al., 2023). We compared the performance of the LoRA implementation in this paper to that of QLoRA (Dettmers et al., 2023), which produced marginally worse results. However, it is possible that other fine-tuning techniques would have produced better results. Further to this, it is possible that if a better model than Llama-3 8B had been fine-tuned we would have seen even better results from the generative LLMs. The computational constraints placed on us are not dissimilar to those that other researchers face, which makes the results in this paper valid while perhaps not exhaustive.
Appendix B Greedy Decoding
Generative models can produce unpredictable outputs (Fadaee and Monz, 2020; Stureborg et al., 2024), which necessitates the use of constrained generation when an LLM forms part of a larger architecture. For the purposes of this work, we use a greedy search to infer the probability of one of the following labels appearing next in the sequence: ”up”, ”down” or ”same”.
Appendix C Dataset Size
@@
Table 4. Dataset sizes
Appendix D Fundamental Data
@@
Table 5. Description of Variables
Appendix E Neural Network Implementations
We also tested some more complex neural network (NN) approaches, which had underwhelming results. The Hierarchical Credit Rating (HierCR) model is a framework that models the filings hierarchically. The challenge with using language models to encode the filings is the limited context window of encoder-based language models. There have been many different solutions to this problem, including sparse attention mechanisms (Beltagy et al., 2020), chunking (Sawhney et al., 2020), and feature-based extraction like the methods above (Loughran and McDonald, 2011; Drinkall et al., 2022). Our NN solution to this problem is to split the filing up into sentences and pass the sentence embeddings through an all-mpnet-base-v2 encoder to produce embeddings for the textual data. The text encoder replicates the structure in (Sawhney et al., 2020), the only material difference is that filing sentences substitute the social media posts in the first layer of the text-encoder. The text, macro and fundamental vectors across the previous quarter(s) are combined using a GRU layer (Chung et al., 2014), the outputs are then passed through an attention layer to create a representation for each data-type. These representations are combined using a bilinear transformation, which is passed through 3 linear layers followed by a ReLU (Agarap, 2019) activation function. Dropout is applied in the final linear-layer classification module.
@@ Appendix E Neural Network Implementations
Table 6. Accuracy for more complex NN approaches using all data types (M+F+T).
The other two architectures are Shared Attention Late Fusion (SA-LF) and Shared Attention Early Fusion (SA-EF). Both architectures only consider the first 512 tokens of each filing. The difference between the two architectures is when the attention layer is applied. For the SA-EF the model attends to all feature-types together, whereas the SA-LF only combines the representations after the attention layer is applied to the individual data-types. The final representation is the passed through the same linear-layer classification module as the HierCR. For all of the architectures above, we trained the model for 200 epochs with a patience value of 20 epochs.
@@ Appendix F Prompts
Your task is to classify the company into one of the following classes: ”down”, ”same”, ”up”. ”down” means that you think the credit rating will go down in the next quarter, meaning the company is perceived as more risky. ”same” means that you think the credit rating will stay the same in the next quarter. ”up” means that you think the credit rating will go up in the next quarter, meaning the company is perceived as less risky. Please respond with a single label that you think fits the company best.
Classify the following numerical data:”””
F.1. Credit Rating Ranking
One potential problem with the prompt outlined in Appendix F is that the LLM may find it hard to correctly understand the ranking structure of credit ratings, which would limit the ability of an LLM to perform well on this task. To probe the LLMs ability to understand the relative rank of credit ratings we created the following prompt:
”””Two credit ratings will be given, the task is to determine which is higher on the following scale, which is ordered in descending order:
@@ F.1. Credit Rating Ranking
¡¡rating_X¿¿ vs. ¡¡rating_Y¿¿ Answer:”””
@@
Appendix G S&P Credit Rating Definitions
@@
Table 7. S&P Global Ratings Definitions (spg, 2016)

### v2

@@ 6 Interpretability
The most obvious advantage of a feature-based system is that important features can be identified. Table 2 provides an example of feature importance that can be used to infer the modality preference of model configurations. It is possible to conduct even more granular feature analysis by looking at the contribution of individual features. Figure 2 shows the partial dependence plot of some of the individual text features on the "Up" and "Down" classes. From the plot in Figure 2(a) we can infer that as ratings are discussed more in a company’s filing, there is a reduced chance of the the credit rating being upgraded. We can also infer from Figure 2(b) that as companies talk about receivables - the money owed to the company - there is a reduced chance of the credit rating being downgraded. Both provide valuable insights and are examples of how a traditional feature-based methodology can be leveraged for increased interpretability.
(a) PDP of "rating_sp_ratings_moodys" cluster & "Up" class.
(b) PDP of "receivables_percent_owned_offs" cluster & "Down" class.
Figure 2: Partial Dependence Plots (PDP) of text-based features against different target classes.
7 Conclusion
@@
The paper shows that while LLMs are good at encoding textual data and inferring signals that traditional methods cannot pick up on when combined with numerical data in the prompt there is performance deterioration. The other advantage is that traditional methods offer increased interpretability and a better understanding of the mechanisms behind certain predictions. In addition, traditional approaches don’t suffer to the same extent from complications associated with training data contamination and memorization Ozdayi et al. (2023); Lu et al. (2024) since the models used for the traditional features are much smaller than the generative models and memorisation in LLMs exhibits scaling law behaviour. While it is not impossible that the text data used in this paper was included in the training of the studied LLMs, any potential influence would likely have a greater impact on generative models, thereby reinforcing the findings of this paper.
@@
There has been some work jointly encoding text using generative LLMs with time-series information Liu et al. (2024), but more work needs to be done to determine the best methodology for combining long text sequences with numerical information while utilizing the benefits of generative LLM natural language understanding. This paper shows that combining multimodal information within the prompt is not sufficient.
8 Limitations
@@
The task above uses a balanced dataset, which is good for testing the different methodologies’ ability to discern the signals that are predictive of a rise or fall in credit ratings, but is poor for assessing how good the models would be in a real-world context where almost all of the ratings stay the same. Despite the data being taken from across all US equities over a 23-year time period, the balanced dataset is relatively small, with only 3441 samples for the Lag 1 configuration and 2142 samples for the Lag 4 configuration. There are plenty of prominent datasets that are smaller, but the size reduces the scope for complex and specialized models to be deployed on this task in favour of more robust, simple models.
@@
Another limitation is that the text used in this paper is produced by the companies themselves, who the goal of conveying a positive viewpoint to investors. More objective publication venues may produce different insights about the future direction of a company.
@@
The LoRA fine-tuning methodology outlined in this paper is a parameter-efficient technique and has been shown to be competitive in a variety of settings Hu et al. (2021), but can be outperformed in some tasks by full fine-tuning and other adapter-based methods Xu et al. (2023). We compared the performance of the LoRA implementation in this paper to that of QLoRA Dettmers et al. (2023), which produced marginally worse results. However, it is possible that other fine-tuning techniques would have produced better results. Further to this, it is possible that if a better model than Llama-3 8B had been fine-tuned we would have seen even better results from the generative LLMs. The computational constraints placed on us are not dissimilar to those that other researchers face, which makes the results in this paper valid while perhaps not exhaustive.
Appendix A Greedy Decoding
Generative models can produce unpredictable outputs Fadaee and Monz (2020); Stureborg et al. (2024), which necessitates the use of constrained generation when an LLM forms part of a larger architecture. For the purposes of this work, we use a greedy search to infer the probability of one of the following labels appearing next in the sequence: "up", "down" or "same".
Appendix B Dataset Size
@@
Table 4: Dataset sizes
Appendix C Fundamental Data
@@
Table 5: Description of Variables
Appendix D Neural Network Implementations
We also tested some more complex neural network (NN) approaches, which had underwhelming results. The Hierarchical Credit Rating (HierCR) model is a framework that models the filings hierarchically. The challenge with using LLMs to encode the filings is the limited context window of encoder-based LLMs. There have been many different solutions to this problem, including sparse attention mechanisms Beltagy et al. (2020), chunking Sawhney et al. (2020), and feature-based extraction like the methods above Loughran and McDonald (2011); Drinkall et al. (2022). Our NN solution to this problem is to split the filing up into sentences and pass the sentence embeddings through an all-mpnet-base-v2 encoder to produce embeddings for the textual data. The text encoder replicates the structure in Sawhney et al. (2020), the only material difference is that filing sentences substitute the social media posts in the first layer of the text-encoder. The text, macro and fundamental vectors across the previous quarter(s) are combined using a GRU layer Chung et al. (2014), the outputs are then passed through an attention layer to create a representation for each data type. These representations are combined using a bilinear transformation, which is passed through 3 linear layers followed by a ReLU Agarap (2019) activation function. Dropout is applied in the final linear-layer classification module.
@@ Appendix D Neural Network Implementations
Table 6: Accuracy for more complex NN approaches using all data types (M+F+T).
The other two architectures are Shared Attention Late Fusion (SA-LF) and Shared Attention Early Fusion (SA-EF). Both architectures only consider the first 512 tokens of each filing. The difference between the two architectures is when the attention layer is applied. For the SA-EF the model attends to all feature types together, whereas the SA-LF only combines the representations after the attention layer is applied to the individual data types. The final representation is the passed through the same linear-layer classification module as the HierCR. For all of the architectures above, we trained the model for 200 epochs with a patience value of 20 epochs.
@@ Appendix E Prompts
Your task is to classify the company into one of the following classes: "down", "same", "up". "down" means that you think the credit rating will go down in the next quarter, meaning the company is perceived as more risky. "same" means that you think the credit rating will stay the same in the next quarter. "up" means that you think the credit rating will go up in the next quarter, meaning the company is perceived as less risky. Please respond with a single label that you think fits the company best.
Classify the following numerical data:"""
E.1 Credit Rating Ranking
One potential problem with the prompt outlined in Appendix E is that the LLM may find it hard to correctly understand the ranking structure of credit ratings, which would limit the ability of an LLM to perform well on this task. To probe the LLMs ability to understand the relative rank of credit ratings we created the following prompt:
"""Two credit ratings will be given, the task is to determine which is higher on the following scale, which is ordered in descending order:
@@ E.1 Credit Rating Ranking
<<rating_X>> vs. <<rating_Y>> Answer:"""
@@
Appendix F S&P Credit Rating Definitions
@@
Table 7: S&P Global Ratings Definitions S&P-Global (2016)